# One-Shot Strategic Classification Under Unknown Costs

### v2

@@ Abstract
The goal of strategic classification is to learn decision rules which are robust to strategic input manipulation. Earlier works assume that these responses are known; while some recent works handle unknown responses, they exclusively study online settings with repeated model deployments. But there are many domains—particularly in public policy, a common motivating use case—where multiple deployments are infeasible, or where even one bad round is unacceptable. To address this gap, we initiate the formal study of one-shot strategic classification under unknown responses, which requires committing to a single classifier once. Focusing on uncertainty in the users’ cost function, we begin by proving that for a broad class of costs, even a small mis-estimation of the true cost can entail trivial accuracy in the worst case. In light of this, we frame the task as a minimax problem, with the goal of identifying the classifier with the smallest worst-case risk over an uncertainty set of possible costs. We design efficient algorithms for both the full-batch and stochastic settings, which we prove converge (offline) to the minimax solution at the dimension-independent rate of $\tilde{\mathcal{O}}(T^{-\nicefrac{{1}}{{2}}})$. Our theoretical analysis reveals important structure stemming from strategic responses, particularly the value of dual norm regularization with respect to the cost function.
@@ 1 Introduction
Motivated by such examples, this work studies robust strategic learning in a “one-shot” setting where the learner must commit to one single model at deployment. In this setting, we recast our objective as optimizing for a classifier that minimizes the worst-case strategic risk over an uncertainty set ${\mathcal{C}}$. We devise two efficient algorithms, for the full-batch and stochastic settings, which provably converge to the minimax solution at the rate $\tilde{\mathcal{O}}(T^{-\nicefrac{{1}}{{2}}})$. Notably, this rate is dimension independent for the typical cost functions—including $\ell_{1}$ and $\ell_{2}$ norms—and it is achieved despite the inner maximization being non-concave. A key step in our approach is to adapt the strategic hinge loss (Levanon and Rosenfeld, 2022) to properly handle a large set of possible costs beyond the $\ell_{2}$ norm. As the strategic hinge loss is non-convex, we apply a regularization term that accounts for the unique structure of strategic response: our analysis uncovers that the “correct” form of regularization is in fact the dual norm of $\beta$ with respect to the cost. We complement this with an updated generalization bound for this loss which corrects a previous error.
@@ 2 Preliminaries
We study the problem of linear strategic classification on inputs $x\in{\mathcal{X}}\subseteq\mathbb{R}^{d}$ and labels $y\in{\mathcal{Y}}:=\{\pm 1\}$ where the exact response of the test population is unknown. We consider linear classifiers $\operatorname{sign}(\beta^{\top}x)$, optimized over some bounded set ${\mathcal{B}}\subset\mathbb{R}^{d+1}$ (this includes a bias term which we leave implicit and which is not included in the vector norm). To model users’ responses, we consider the typical setup of a cost function $c(x,x^{\prime})$ which defines the cost for an agent to change their features from $x$ to $x^{\prime}$. Together with a utility $u\geq 0$ gained from a positive classification, this cost determines the strategic response of a rational agent to a classifier $\beta$ via
@@
Recall that we are interested in robust prediction when we cannot know the exact strategic response. Keeping the assumption of rational behavior, this naturally points to an unknown cost function as a primary source of this uncertainty. We parameterize this uncertainty via an unknown positive definite (PD) matrix $\Sigma\succ\mathbf{0}$ in $\mathbb{R}^{d\times d}$ which scales the relative costs per input dimension, denoting the induced norm as $\|\cdot\|_{\Sigma}$. For the 2-norm, this is the standard PD norm given by $\|x\|_{\Sigma}:=\sqrt{x^{\top}\Sigma x}$, but we define it for general norms as $\|x\|_{\Sigma}:=\|\Sigma^{1/2}x\|$. Hence, for our purposes, any cost function $c$ is uniquely determined by its parameterization $\Sigma$; we will use these two variables interchangeably. The other two factors in strategic response are the positive utility $u$ and the monotone transform $\phi$: for reasons that will soon be made clear, we only require knowledge of the maximum value $u_{*}$ such that $\phi(u_{*})\leq u$ (this is clearly satisfied for known $\phi$ and $u$, as is typically assumed). Note this value need not be shared among users, and it suffices to know an interval in which it lies—but for simplicity we treat it as fixed.
@@ 2.1 Strategic Learning Under a Single, Known Cost
Unfortunately, even this relaxation poses difficulties. Firstly, the objective is non-convex—though Levanon and Rosenfeld (2022) show that for known costs it often learns reasonable classifiers in practice, once we transition to unknown costs it will become clear that having a guaranteed sub-optimality bound is important (such a bound is impossible in general for non-convex objectives). Second, the additional term $2\|\beta\|_{2}$ captures the effect of strategic response only under the standard $\ell_{2}$-norm cost. [^1] While Levanon and Rosenfeld (2022) do discuss more general costs, they give only a generic description.
@@
The $\Sigma$-transformed dual norm of $\beta$ is denoted $\|\beta\|_{*,\Sigma^{-1}}:=\sup_{\|v\|_{\Sigma}=1}\beta^{\top}v$ and is equal to $\|\Sigma^{-\nicefrac{{1}}{{2}}}\beta\|_{*}$. We may leave dependence on $\Sigma$ implicit, writing simply $\|\beta\|_{*}$.
@@
Fix $\beta$ and let $c(x,x^{\prime})=\phi(\|x-x^{\prime}\|_{\Sigma})$. The maximum possible change to a user’s score due to strategic behavior is $u_{*}\|\beta\|_{*}$.
@@
Denote the overall strategic hinge risk by $R_{\textrm{s-hinge}}^{c}(\beta):=\mathbb{E}[\ell_{\textrm{s-hinge}}^{c}(\beta^{\top}x,y)]$, with the strategic 0-1 risk defined analogously as $R_{0-1}^{c}$. By design, the former upper bounds the latter:
@@ 3 The Perils of Using a Wrong Cost Function
Suppose one instead learns a classifier $\hat{\beta}$ by assuming an incorrect cost $\hat{\Sigma}$ and minimizing the population strategic 0-1 risk under that cost: $\hat{\beta}=\operatorname*{arg\,min}_{\beta}\mathbb{E}_{q}[\ell_{0-1}^{\hat{c}}(\beta)]$. Then the excess 0-1 risk suffered by $\hat{\beta}$ is
@@
where $\epsilon=\epsilon(\Sigma,\hat{\Sigma},\mu_{0}):=\frac{u_{*}|\|\mu_{0}\|_{*,\hat{\Sigma}^{-1}}-\|\mu_{0}\|_{*,\Sigma^{-1}}|}{\|\mu_{0}\|}$ is the normalized estimation error of the $\Sigma$-transformed dual norm of $\mu_{0}$.
Theorem 3.2 demonstrates that even for much more benign distributions, choosing a classifier based on a slightly incorrect cost can be very non-robust. Roughly, we should expect $\epsilon=\Theta(\textrm{Trace}(\hat{\Sigma}^{-1}-\Sigma^{-1}))$, scaling as our error grows along directions where $\mu_{0}$ is large. Crucially, this is with respect to the inverse cost matrix $\Sigma^{-1}$, which means that a tiny estimation error in the lower end of the eigen-spectrum can have a large effect. For example, imagine that the classes are sufficiently separated so that the optimal classifier absent strategic behavior can get accuracy close to 1 (which is trivial in high dimensions). Then under strategic behavior, the error in guessing the smallest eigenvalue of $\Sigma$ need only be $\Omega(\|\mu_{0}\|\lambda_{\min}(\Sigma)\lambda_{\min}(\hat{\Sigma}))$ to induce excess risk of approximately $\nicefrac{{1}}{{4}}$, and it will rapidly approach $\nicefrac{{1}}{{2}}$ as $d$ grows. To give a bit more intuition for these factors, Figure 1 visualizes a few simple examples of how the excess risk behaves as dimension increases, depending on the distribution of eigenvalues of $\Sigma$ and our error in estimating them. More abstractly, we can see that if we only slightly err in estimating the cost of movement in a direction, it could cause a very large error in estimating how much agents will move in that direction.
Figure 1: Visualizing the excess 0-1 risk from Theorem 3.2. Left: A toy illustration of the excess risk as a function of $\|\mu_{0}\|$ and $\epsilon$ and where they lie on the Gaussian CDF. Right: Excess risk curves as a function of the dimension $d$, where $\mu_{0}=\frac{1}{\sqrt{d}}\mathbf{1}$ and $\sigma^{2}=\nicefrac{{1}}{{d}}$. Color indicates the function which generates the spectrum of $\Sigma$, where $\lambda_{k}$ is the $k$-largest eigenvalue in the series. Line style indicates the error $\varepsilon$ in estimating the eigenvalue in each dimension—precisely, we define $\hat{\Sigma}=(1-\varepsilon)\Sigma$. Even with all eigenvalues estimated to error $1-e^{-10}$, excess risk grows rapidly with $d$ towards $\frac{1}{2}$.
@@ 4.1 Solving the Inner Max for a Fixed Classifier
     Otherwise, return $\Sigma\in{\mathcal{C}}$ such that $\|\beta\|_{*,\Sigma^{-1}}=k_{\max}$.
@@ 4.2 Solving the Outer Min for a Fixed Cost
Proofs are in Appendix B. Proposition 4.3 shows that a small, cost-independent amount of dual norm regularization ensures convexity of the outer min problem for any cost function, making it the natural choice in this setting.
@@ 5 Efficiently Identifying the Minimax-Optimal Classifier
$\min_{\beta}\max_{{c\in{\mathcal{C}}}}\bigl{[}R_{\textrm{s-hinge}}^{c}(\beta)+\lambda u_{*}\|\beta\|_{*}\bigr{]}.$ (4)
@@ 5.1 Solving the Objective in the Full-Batch Setting with the Subgradient Method
$\max_{{c\in{\mathcal{C}}}}R_{0-1}^{c}(\hat{\beta})\leq$ $\min_{\beta}\max_{{c\in{\mathcal{C}}}}R_{\textrm{s-hinge}}^{c}(\beta)+\mathcal{O}\left(\frac{LB}{\sqrt{T}}+(X+u_{*})\frac{B+\sqrt{\ln\nicefrac{{1}}{{\delta}}}}{\sqrt{n}}\right).$
@@ 5.2 Solving the Objective in the Minibatch Setting with Stochastic Mirror Descent-Ascent
Modifying mirror ascent to apply to our setting.
where $\Delta(|\mathcal{S}|)$ is the $|\mathcal{S}|$-simplex. In other words, we can solve this problem over all convex combinations of the risks under the different costs in $\mathcal{S}$ and arrive at the same solution (Rosenfeld et al., 2022). Crucially, unlike our original objective, Equation 5 is convex-concave, so it can be efficiently optimized via SMDA! The correctness of this relaxation relies on the property proven in Lemma 2.3: strictly speaking, the strategic loss depends only the scalar dual norm, not on the full cost matrix.
@@
$\mathbb{E}[\max_{{c\in{\mathcal{C}}}}R_{\textrm{s-hinge}}^{c}(\hat{\beta})]$ $\lesssim\min_{\beta}\max_{{c\in{\mathcal{C}}}}R_{\textrm{s-hinge}}^{c}(\beta)+\mathcal{O}\left(\frac{LB}{\sqrt{T}}+B(X+u_{*})\sqrt{\frac{\ln TD}{T}}\right).$
@@ Appendix A Proofs of Hardness Results
where $\epsilon:=\frac{u_{*}|\|\mu_{0}\|_{*,\hat{\Sigma}^{-1}}-\|\mu_{0}\|_{*,\Sigma^{-1}}|}{\|\mu_{0}\|}$.
@@
By the Neyman-Pearson Lemma, the classifier which minimizes non-strategic 0-1 risk will be the one which predicts $\operatorname{sign}(\ln\frac{p(y=1\mid x)}{p(y=0\mid x)})$, which gives $\beta^{*}=2\mu_{0}$. To account for strategic response, we observe that as proven in Lemma 2.3, each user will have $\delta(x)$ in the direction of $\beta^{*}$ which induces a change in their predicted value by at most $u_{*}\|\beta^{*}\|_{*}=2u_{*}\|\mu_{0}\|_{*,\Sigma^{-1}}$; since we are considering the 0-1 loss, this is equivalent to every user shifting in this way.
It is immediate that to find the corresponding minimizer of the population 0-1 strategic risk, we can simply add a negative bias equal to the change induced by this shift, because this maintains the same labels as before on all points, after they strategically shift. Therefore, the strategic risk minimizer will be $2\mu_{0}^{\top}x-2u_{*}\|\mu_{0}\|_{*,\Sigma^{-1}}$. By the same argument, the minimizer of the 0-1 strategic risk under the incorrect cost matrix $\hat{\Sigma}$ is $2\mu_{0}^{\top}x-2u_{*}\|\mu_{0}\|_{*,\hat{\Sigma}^{-1}}$. It remains to derive a lower bound on their difference in risk.
First note that as argued above, the 0-1 strategic risk of the minimizer for the correct cost is the same as the non-strategic risk of the non-strategic solution. For the both positively and negatively labeled points, this is equal to $\Phi\left(-\frac{\|\mu_{0}\|}{\sigma}\right)$ due to rotational symmetry. To determine the risk of the incorrect solution, we can identify the regions whose label will differ under that classifier and bound the measure of those regions. Define $\gamma=2u_{*}(\|\mu_{0}\|_{*,\hat{\Sigma}^{-1}}-\|\mu_{0}\|_{*,\Sigma^{-1}})$ as the difference in the two solutions’ predictions on all $x$. Due to the symmetry of the positive and negative conditional distributions, we can assume WLOG that $\gamma>0$, i.e., the incorrect classifier assigns a smaller prediction to all $x$. This means it will have less risk on a region of negative points and more on positive. Specifically, the two classifiers will differ on all points for which the true strategic-optimal classifier assigns a value before strategic response which lies in $(-2u_{*}\|\mu_{0}\|_{*,\Sigma^{-1}},-2u_{*}\|\mu_{0}\|_{*,\Sigma^{-1}}+\gamma)$; those assigned a value less than $-2u_{*}\|\mu_{0}\|_{*,\Sigma^{-1}}$ will receive a negative prediction from both classifiers, and those assigned a value greater than $-2u_{*}\|\mu_{0}\|_{*,\Sigma^{-1}}+\gamma$ will still be close enough to the decision boundary that they can shift to achieve a positive label prediction from $\hat{\beta}$. Formally, this region is
$\{x\mid-2u_{*}\|\mu_{0}\|_{*,\Sigma^{-1}}<\beta^{*\top}x<-2u_{*}\|\mu_{0}\|_{*,\Sigma^{-1}}+\gamma\}$ =$=\{x\mid 0<\mu_{0}^{\top}x<\nicefrac{{\gamma}}{{2}}\}.$ (14)
@@
where $\epsilon:=\frac{u_{*}|\|\mu_{0}\|_{*,\hat{\Sigma}^{-1}}-\|\mu_{0}\|_{*,\Sigma^{-1}}|}{\|\mu_{0}\|}$. By applying the fundamental theorem of calculus and the fact that $\Phi(x)=1-\Phi(-x)$ we arrive at the stated equality. ∎

### v3

@@ Abstract
The goal of strategic classification is to learn decision rules which are robust to strategic input manipulation. Earlier works assume that these responses are known; while some recent works handle unknown responses, they exclusively study online settings with repeated model deployments. But there are many domains—particularly in public policy, a common motivating use case—where multiple deployments are infeasible, or where even one bad round is unacceptable. To address this gap, we initiate the formal study of one-shot strategic classification under unknown responses, which requires committing to a single classifier once. Focusing on uncertainty in the users’ cost function, we begin by proving that for a broad class of costs, even a small mis-estimation of the true cost can entail trivial accuracy in the worst case. In light of this, we frame the task as a minimax problem, aiming to minimize worst-case risk over an uncertainty set of costs. We design efficient algorithms for both the full-batch and stochastic settings, which we prove converge (offline) to the minimax solution at the rate of $\tilde{\mathcal{O}}(T^{-\nicefrac{{1}}{{2}}})$. Our analysis reveals important structure stemming from strategic responses, particularly the value of dual norm regularization with respect to the cost function.
Machine Learning, ICML
@@ 1 Introduction
Motivated by such examples, this work studies robust strategic learning in a “one-shot” setting where the learner must commit to one single model at deployment. In this setting, we recast our objective as optimizing for a classifier that minimizes the worst-case strategic risk over an uncertainty set ${\mathcal{C}}$. We devise two efficient algorithms, for the full-batch and stochastic settings, which provably converge to the minimax solution at the rate $\tilde{\mathcal{O}}(T^{-\nicefrac{{1}}{{2}}})$. Notably, this rate is dimension independent for the typical cost functions—including $\ell_{1}$ and $\ell_{2}$ norms—and it is achieved despite the inner maximization being non-concave. A key step in our approach is to adapt the strategic hinge loss (Levanon and Rosenfeld, 2022) to properly handle a large set of possible costs beyond the $\ell_{2}$ norm. As the strategic hinge loss is non-convex, we apply a regularization term that accounts for the unique structure of strategic response: our analysis uncovers that the “correct” form of regularization is in fact the dual norm of $\beta$ with respect to the cost. We complement this with an updated generalization bound for this loss which corrects an error in that previous work.
@@ 2 Preliminaries
We study the problem of linear strategic classification on a distribution over inputs $x\in{\mathcal{X}}\subseteq\mathbb{R}^{d}$ and labels $y\in{\mathcal{Y}}=\{\pm 1\}$ where the exact response of the test population is unknown. We consider linear classifiers $\operatorname{sign}(\beta^{\top}x)$, optimized over some bounded set ${\mathcal{B}}\subset\mathbb{R}^{d+1}$ (this includes a bias term which we leave implicit and which is not included in the vector norm). To model users’ responses, we consider the typical setup of a cost function $c(x,x^{\prime})$ which defines the cost for an agent to change their features from $x$ to $x^{\prime}$. Together with a utility $u\geq 0$ gained from a positive classification, this cost determines the strategic response of a rational agent to a classifier $\beta$ via
@@
Recall that we are interested in robust prediction when we cannot know the exact strategic response. Keeping the assumption of rational behavior, this naturally points to an unknown cost function as a primary source of this uncertainty. We parameterize this uncertainty via an unknown positive definite (PD) matrix $\Sigma\succ\mathbf{0}$ in $\mathbb{R}^{d\times d}$ which scales the relative costs per input dimension, denoting the induced norm as $\|\cdot\|_{\Sigma}$. For the 2-norm, this is the standard PD norm given by $\|x\|_{\Sigma}:=\sqrt{x^{\top}\Sigma x}$, but we define it for general norms as $\|x\|_{\Sigma}:=\|\Sigma^{1/2}x\|$. Hence, any cost function $c$ is uniquely determined by its parameterization $\Sigma$ as $c(x,x^{\prime})=\phi\left(\|\Sigma^{1/2}(x^{\prime}-x)\|\right)$; we will use these two variables interchangeably. The other two factors in strategic response are the positive utility $u$ and the monotone transform $\phi$: for reasons that will soon be made clear, we only require knowledge of the maximum value $u_{*}$ such that $\phi(u_{*})\leq u$ (this is clearly satisfied for known $\phi$ and $u$, as is typically assumed). Note this value need not be shared among users, and it suffices to know an interval in which it lies—but for simplicity we treat it as fixed.
@@ 2.1 Strategic Learning Under a Single, Known Cost
Unfortunately, even this relaxation poses difficulties. Firstly, the objective is non-convex—though Levanon and Rosenfeld (2022) show that for known costs it often learns reasonable classifiers in practice, once we transition to unknown costs it will become clear that having a guaranteed sub-optimality bound, which requires convexity, is important. Second, the additional term $2\|\beta\|_{2}$ captures the effect of strategic response only under the standard $\ell_{2}$-norm cost. [^1] While Levanon and Rosenfeld (2022) do discuss more general costs, they give only a generic description.
@@
The $\Sigma$-transformed dual norm of $\beta$ is denoted $\|\beta\|_{*,\Sigma}:=\sup_{\|v\|_{\Sigma}=1}\beta^{\top}v=\|\Sigma^{-\nicefrac{{1}}{{2}}}\beta\|_{*}$. We may leave dependence on $\Sigma$ implicit, writing simply $\|\beta\|_{*}$.
@@
Fix $\beta$ and let $c(x,x^{\prime})=\phi(\|x-x^{\prime}\|_{\Sigma})$. The maximum possible change to a user’s score due to strategic behavior is $u_{*}\|\beta\|_{*,\Sigma}$.
@@
Denote the overall strategic hinge risk by $R_{\textrm{s-hinge}}^{c}(\beta):=\mathbb{E}[\ell_{\textrm{s-hinge}}^{c}(\beta^{\top}x,y)]$, with the strategic 0-1 risk defined analogously as $R_{0-1}^{c}$.
@@ 3 The Perils of Using a Wrong Cost Function
where $\epsilon:=\frac{u_{*}|\|\mu_{0}\|_{*,\hat{\Sigma}}-\|\mu_{0}\|_{*,\Sigma}|}{\|\mu_{0}\|}$.
Theorem 3.2 demonstrates that even for much more benign distributions, choosing a classifier based on a slightly incorrect cost can be very non-robust. Roughly, we should expect $\epsilon=\Theta(u_{*}\textrm{tr}(\hat{\Sigma}^{-1}-\Sigma^{-1}))$, scaling as our error grows along directions where $\mu_{0}$ is large. Crucially, this is with respect to the inverse cost matrix $\Sigma^{-1}$, which means that a tiny estimation error in the lower end of the eigen-spectrum can have a large effect. For example, imagine that the classes are sufficiently separated so that the optimal classifier absent strategic behavior can get accuracy close to 1 (which is trivial in high dimensions). Then under strategic behavior, the error in guessing the smallest eigenvalue of $\Sigma$ need only be $\Omega(\|\mu_{0}\|\lambda_{\min}(\Sigma)\lambda_{\min}(\hat{\Sigma}))$ to induce excess risk of approximately $\nicefrac{{1}}{{4}}$, and it will rapidly approach $\nicefrac{{1}}{{2}}$ as $d$ grows. To give a bit more intuition for these factors, Figure 1 visualizes a few simple examples of how the excess risk behaves as dimension increases, depending on the distribution of eigenvalues of $\Sigma$ and our error in estimating them. More abstractly, we can see that if we only slightly err in estimating the cost of movement in a direction, it could cause a very large error in estimating how much agents will move in that direction.
@@ 4.1 Solving the Inner Max for a Fixed Classifier
     Otherwise, return $\Sigma\in{\mathcal{C}}$ such that $\|\beta\|_{*,\Sigma}=k_{\max}$.
@@ 4.2 Solving the Outer Min for a Fixed Cost
In fact, we can show that $q(y=1)u_{*}\|\beta\|_{*}$ is the minimum regularization that ensures convexity: for any $\lambda$, if the objective in Theorem 4.2 is convex, then $q(y=1)\|\beta\|_{*}\leq\lambda\|\beta\|_{2}$. Proofs are in Appendix B. Proposition 4.3 shows that a small, cost-independent amount of dual norm regularization ensures convexity of the outer min problem for any cost function, making it the natural choice in this setting.
@@ 5 Efficiently Identifying the Minimax-Optimal Classifier
$\min_{\beta}\max_{{c\in{\mathcal{C}}}}\bigl{[}R_{\textrm{s-hinge}}^{c}(\beta)+\lambda u_{*}\|\beta\|_{*,\Sigma}\bigr{]}.$ (4)
@@ 5.1 Solving the Objective in the Full-Batch Setting with the Subgradient Method
$\max_{{c\in{\mathcal{C}}}}R_{0-1}^{c}(\hat{\beta})\leq$ $\min_{\beta}\max_{{c\in{\mathcal{C}}}}R_{\textrm{s-hinge}}^{c}(\beta)$
$+\mathcal{O}\left(\frac{LB}{\sqrt{T}}+(X+u_{*})\frac{B+\sqrt{\ln\nicefrac{{1}}{{\delta}}}}{\sqrt{n}}\right).$
@@ 5.2 Solving the Objective in the Minibatch Setting with Stochastic Mirror Descent-Ascent
Modifying mirror ascent to apply to our setting.
where $\Delta(|\mathcal{S}|)$ is the $|\mathcal{S}|$-simplex. In other words, we can solve this problem over all convex combinations of the risks under the different costs in $\mathcal{S}$ and arrive at the same solution (Rosenfeld et al., 2022a, b). Crucially, unlike our original objective, Equation 5 is convex-concave, so it can be efficiently optimized via SMDA! The correctness of this relaxation relies on the property proven in Lemma 2.3: strictly speaking, the strategic loss depends only the scalar dual norm, not on the full cost matrix.
@@
$\mathbb{E}[\max_{{c\in{\mathcal{C}}}}R_{\textrm{s-hinge}}^{c}(\hat{\beta})]$ $\lesssim\min_{\beta}\max_{{c\in{\mathcal{C}}}}R_{\textrm{s-hinge}}^{c}(\beta)$
$+\mathcal{O}\left(\frac{LB}{\sqrt{T}}+B(X+u_{*})\sqrt{\frac{\ln TD}{T}}\right).$
@@ 6 Conclusion
Impact Statement
This work studies the problem of strategic classification under unknown user costs. While our focus is entirely theoretical, it is important to acknowledge that, in a realistic setting, the learning task which we consider would have human beings as the targets for prediction. This is in fact our primary motivation: correctly accounting for strategic behavior in learning can benefit not only the system, but also the general population of users (Milli et al., 2019; Levanon and Rosenfeld, 2021, 2022). Questions of how to act under uncertainty or partial information are immanent in strategic learning; our work offers but one way of accounting for system uncertainty regarding user behavior, with a particular focus on uncertainty stemming from unknown user costs. Other sources of uncertainty therefore lie outside our scope, as does human behavior which deviates from our assumed rational response model (Eq. equation 1), as well as considerations beyond minimizing worst case risk. Although our analysis applies to a relatively large and flexible class of cost functions (broader than what is typically considered in the literature), a clear limitation of our work is that it is unclear if it applies to costs outside this class. A more subtle but nonetheless important limitation is that while our approach provides strong worst-case guarantees for any cost in the uncertainty set ${\mathcal{C}}$, it does not state what will happen if the ‘true’ cost $c^{*}$ lies outside of the chosen ${\mathcal{C}}$. We conjecture that under appropriate smoothness assumptions, worst-case guarantees should degrade gracefully as $c^{*}$ moves away from ${\mathcal{C}}$, but leave this as an open question for future work. Finally, although the algorithms we propose are theoretical, they may nonetheless be applicable in certain settings. Given this, we implore any future use of our algorithms to be carried out in a transparent and accountable manner—as should be expected from any tool developed under the general framework of strategic learning.
@@ Appendix A Proofs of Hardness Results
where $\epsilon:=\frac{u_{*}|\|\mu_{0}\|_{*,\hat{\Sigma}}-\|\mu_{0}\|_{*,\Sigma}|}{\|\mu_{0}\|}$.
@@
By the Neyman-Pearson Lemma, the classifier which minimizes non-strategic 0-1 risk will be the one which predicts $\operatorname{sign}(\ln\frac{p(y=1\mid x)}{p(y=0\mid x)})$, which gives $\beta^{*}=2\mu_{0}$. To account for strategic response, we observe that as proven in Lemma 2.3, each user will have $\delta(x)$ in the direction of $\beta^{*}$ which induces a change in their predicted value by at most $u_{*}\|\beta^{*}\|_{*}=2u_{*}\|\mu_{0}\|_{*,\Sigma}$; since we are considering the 0-1 loss, this is equivalent to every user shifting in this way.
It is immediate that to find the corresponding minimizer of the population 0-1 strategic risk, we can simply add a negative bias equal to the change induced by this shift, because this maintains the same labels as before on all points, after they strategically shift. Therefore, the strategic risk minimizer will be $2\mu_{0}^{\top}x-2u_{*}\|\mu_{0}\|_{*,\Sigma}$. By the same argument, the minimizer of the 0-1 strategic risk under the incorrect cost matrix $\hat{\Sigma}$ is $2\mu_{0}^{\top}x-2u_{*}\|\mu_{0}\|_{*,\hat{\Sigma}}$. It remains to derive a lower bound on their difference in risk.
First note that as argued above, the 0-1 strategic risk of the minimizer for the correct cost is the same as the non-strategic risk of the non-strategic solution. For the both positively and negatively labeled points, this is equal to $\Phi\left(-\frac{\|\mu_{0}\|}{\sigma}\right)$ due to rotational symmetry. To determine the risk of the incorrect solution, we can identify the regions whose label will differ under that classifier and bound the measure of those regions. Define $\gamma=2u_{*}(\|\mu_{0}\|_{*,\hat{\Sigma}}-\|\mu_{0}\|_{*,\Sigma})$ as the difference in the two solutions’ predictions on all $x$. Due to the symmetry of the positive and negative conditional distributions, we can assume WLOG that $\gamma>0$, i.e., the incorrect classifier assigns a smaller prediction to all $x$. This means it will have less risk on a region of negative points and more on positive. Specifically, the two classifiers will differ on all points for which the true strategic-optimal classifier assigns a value before strategic response which lies in $(-2u_{*}\|\mu_{0}\|_{*,\Sigma},-2u_{*}\|\mu_{0}\|_{*,\Sigma}+\gamma)$; those assigned a value less than $-2u_{*}\|\mu_{0}\|_{*,\Sigma}$ will receive a negative prediction from both classifiers, and those assigned a value greater than $-2u_{*}\|\mu_{0}\|_{*,\Sigma}+\gamma$ will still be close enough to the decision boundary that they can shift to achieve a positive label prediction from $\hat{\beta}$. Formally, this region is
$\{x\mid-2u_{*}\|\mu_{0}\|_{*,\Sigma}<\beta^{*\top}x<-2u_{*}\|\mu_{0}\|_{*,\Sigma}+\gamma\}$ =$=\{x\mid 0<\mu_{0}^{\top}x<\nicefrac{{\gamma}}{{2}}\}.$ (14)
@@
where $\epsilon:=\frac{u_{*}|\|\mu_{0}\|_{*,\hat{\Sigma}}-\|\mu_{0}\|_{*,\Sigma}|}{\|\mu_{0}\|}$. By applying the fundamental theorem of calculus and the fact that $\Phi(x)=1-\Phi(-x)$ we arrive at the stated equality. ∎
@@
As mentioned in the main body, we can also show the following additional result: for any $\lambda$ which induces convexity with $\|\beta\|_{2}$, it must hold that $\|\beta\|_{*}\leq\|\beta\|_{2}$, and thus it imposes the least regularization.
Proposition B.2.
If the objective $R_{\textrm{s-hinge}}^{c}(\beta)+\lambda u_{*}\|\beta\|_{2}$ is convex, then the dual regularization is no greater than the $\ell_{2}$ regularization: $\tau^{+}\|\beta\|_{*}\leq\lambda\|\beta\|_{2}$.
Proof.
Recall Equation 24 in the proof above: $f$ is convex only if for all vectors $\beta_{1},\beta_{2}\in\mathcal{B}$,
Choosing $\beta_{1}=-\beta_{2}$ and simplifying, we get $\lambda||\beta_{2}||_{2}\geq\tau^{+}||\beta_{2}||_{*}$